

********************************************************************
UARS WEEKLY STATUS REPORT - 03 SEP 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this past
week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing reducing activation time
regarding the HEPS Energy Processor trouble. Updated command loads were
sent to the spacecraft on September 1. Detailed telemetry is now being
examined to ensure that compressing the activation sequence did not effect
the instrument.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
***********
The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII was switched off on September 2, after a normal period
of operation. It will be activated again for September 10-14.
Re-processing with V5 is continuing.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
DAAC data created in the month of June 1999 has been loaded into the
DAAC Transfer Tables for HALOE, PEM, NMC, and UKMO.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
15 days CLA (7.0) PROD (L1-L3) (473R-224C-14F=235L)
2 days GPI (4.05) PROD (CLAES L3B) (191R-164C-2F=25L)
0 days WIN (5.11) PROD (L1-L3) (1540R-1028C-33F=479L)
UARS PCG will load July 1999 current day DAAC data in September 1999.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 August, 1999 (Orbit 43402, GMT Day 232) through 27 August,
1999 (Orbit 43505, GMT Day 239).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 235/1102 43438 On
PEM 238/1549 43486 Calibration
WINDII 239/0007 43492 On
The beta angle ranged between the angles of 1.2 degrees to 26.9 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
51 degrees on 07 September 1999 (DOY 250).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.1 degrees C. Additional battery performance data for each
battery on 20 August 1999 and 27 August 1999 are provided in Appendix D.
The clock error ranged from -14.3 to 10.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 06
% HALOE = 02
% WINDII = 01
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/208: 2287.496910
% Transponder B: 99/209: 2287.498901
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/231 show the observatory data loss to be 24 hours, 55
minutes, 09.204 seconds (an increase of 18.912 seconds since the last report
period). This is a 0.0359 percent data loss which equals a 99.9641 percent
data capture for the mission.
The increase of 18.912 seconds occurred over one (2) days:
% 1.056 seconds on DOY 234 due to data hits
% 17.856 seconds on DOY 236 due to data hits, no overlap
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to investigate the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun pointing
safehold.
Dimitrios Mantziaras and Robert Volgelsang began testing the replaced CNE
connection from OBCSTF and SIM2 to the open Lan. As of Friday the new OBCSTF
connection worked just fine, but the SIM2 did not work properly. The SIM2
machine has been acting up lately, so it has not yet been determined whether it
is an UTTS problem or a connection problem.
A negative clock adjust this week. This was an atypical event which was
executed nominally.
YEAR 2000 OPERABILITY SUPPORT
None to report this period.
MEETINGS
FOT personnel are supporting the IMOC (SCS-21) development effort. A
demonstration was held on Thursday, August 26, 1999 defining the capability of
the system.
FUTURE SIGNIFICANT EVENTS
September, 1999: A meeting is planned for the beginning of September to
discuss the UARS Safehold configuration and concerns for the MPS and its
batteries. This meeting will be a closed meeting with only the FOT and
specialized engineering support attending.
Fix PEMUs timing to execute activation sequence quicker.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 August 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43411 233/1423 99-298 OBC Clk Adj fr 0.00 to 20.6 msec/day
43413 233/1854 99-299 OBC Clk Adj fr 20.6 to 0.00 msec/day
43438 235/1102 99-300 HALOE Power ON
43440 235/1414 99-301 OBC Clk Adj fr 0.00 to -2.0 msec/day
43444 235/2030 99-302 HALOE Power ON part 2
43457 236/1558 99-303 OBC Clk Adj fr -20.6 to 0.00 msec/day
43470 237/1348 99-304 OBC Clk Adj fr 0.00 to 20.6 msec/day
43486 238/1549 99-305 PEM PEM CALIBRATION
43487 238/1729 99-306 OBC Clk Adj fr 20.6 to 0.00 msec/day
43492 239/0007 99-307 WINDII Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 August 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
No Out-of-Limits to report this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 August 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
43411 233/14:28 (21 Aug 99) from 00.0 to 20.6 msec/day
43413 233/19:02 (21 Aug 99) from 20.6 to 00.0 msec/day
43440 235/14:18 (23 Aug 99) from 00.0 to -20.6 msec/day
43457 236/16:04 (24 Aug 99) from -20.6 to 00.0 msec/day
43470 237/13:58 (25 Aug 99) from 00.0 to 20.6 msec/day
43487 238/17:34 (26 Aug 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 August 1999
20 August 1999 - GMT Day 232 Beta = 1.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ----- ------- ------ ------ ------
1 -112.0/-134.4 -3.05/-4.62 +0.0/+0.0 16.5 0.000 0.0
2 +22.4/-33.6 +5.47/+4.19 +29.6/-13.6 24.3 1.012 21.4
3 +67.2/+0.0 +4.51/+2.92 +29.2/-16.0 24.2 1.014 21.1
27 August 1999 - GMT Day 239 Beta = 26.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ----- ------- ------
1 -112.0/-134.4 -2.73/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +67.2/+0.0 +6.44/+4.51 +30.8/-15.2 24.3 1.023 22.0
3 +39.2/-5.6 +5.47/+3.87 +30.4/-16.0 24.2 1.022 21.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ----- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 August 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-041 FOT 43442 235/1554 N/A TAC 6 Crashed
99-042 FOT 43445 235/2212 N/A TAC 6 Crashed
99-043 FOT 43458 236/1855 21265 Late Acq / sched.conflict
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list August 99 14170 Generic Late Acq.
SEVNTFW99/35